OpenStudy (anonymous):

a ball is thrown upward from the top of a building, the function below shows the height of the ball above the ground f(t)in feet at different times (t) in seconds: f(t)=-16t^2+32t+90 the average rate of change of f(x) from x=4 to x=6 is?

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

Isn't it just (f(6) - f(4))/2?

OpenStudy (anonymous):

im not sure @IAmSinged

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

yeah it's just the change in f(x) divided by the change in x. http://tutorial.math.lamar.edu/Classes/CalcI/Tangents_Rates.aspx

OpenStudy (anonymous):

wait so you just dived 4 by 6?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

You're finding the slope between the 2 pts. I know the change in y (f(6) - f(4)), and I know the change of x (6 - 4 = 2). Then I just divide the change in y by the change in x, because that's the definition of the slope.

OpenStudy (anonymous):

im confused :/

OpenStudy (anonymous):

You want to find the slope between the points at t = 4 and t = 6 (definition of average rate of change). You get the y values of the points by plugging it into the equation. At t = 4, we get -38. At t =6, we get -294. So we have the points (4, -38) and (6, -294) and we want to find the slope. So we do: (-294 + 38)/(6 - 4) = -128.

OpenStudy (ranga):

The average rate of change of a function, f(x), between x = a and x = b is defined as: \[\frac{f(b) - f(a)}{b-a}\]

OpenStudy (anonymous):

You can prove that the slope between the two points is the average rate of change by finding the derivative divided by the change in x, and then integrating it.

OpenStudy (anonymous):

@ranga so I would plug In 4 and 6? f(4)-f(6) /4-6

OpenStudy (ranga):

Yes, but I would do it the other way (you will get the same answer): { f(6) - f(4) } / { 6 - 4}

OpenStudy (anonymous):

is the answer 1?

OpenStudy (anonymous):

@ranga ?

OpenStudy (ranga):

No, that is not the answer. f(6) = ? f(4) = ?

OpenStudy (anonymous):

6 and 4 right?

OpenStudy (ranga):

f(t)=-16t^2+32t+90 f(6) = -16(6)^2 + 32(6) + 90 = ? f(4) = -16(4)^2 + 32(4) + 90 = ?

OpenStudy (anonymous):

f(6)=-294 f(4)=-38

OpenStudy (ranga):

Yes. Plug it into \[\frac{f(6) - f(4)}{6-4}\]

OpenStudy (anonymous):

-294 - (-38) /-294 -(-38) =-256/-256?

OpenStudy (ranga):

\[ \frac{f(6) - f(4)}{6-4} = \frac{-294+38}{2} = \frac{-256}{2} = -128 ~ft./sec \]

OpenStudy (anonymous):

ohhhhh